OpenStudy (anonymous):

What is the value of n in the equation 12n + 8 = 11? Please submit your answer as a reduced fraction. Its a grid response please help

OpenStudy (anonymous):

you have to change 8 to the right side so you get 12n=11-8

OpenStudy (anonymous):

12n + 8 =11 - 8 - 8 ------------- 12n = 3 12n = 3 --- - 12 12 n = 3/12 n = 1/4 Hope this helps! :D P.S. What do you mean by grid response?

OpenStudy (anonymous):

so you know 12n=3, what you have to do is pass the 12 that is multiplying to the other side dividing so you get n=3/12

OpenStudy (anonymous):

reducing the fraction you find that n=1/4

OpenStudy (anonymous):

thanks to both of you and Andrew i mean that its not multiple choice

OpenStudy (anonymous):

you're welcome :)